# On the Identification of Temporally Causal Representation with Instantaneous Dependence

### v2

@@ 6 Conclusion
$\begin{gathered}\begin{bmatrix}\begin{array}[]{c}z_{t-1,1}\\ z_{t-1,2}\\ z_{t,1}\\ z_{t,2}\end{array}\end{bmatrix}=\mathbf{f}\left(\begin{bmatrix}\begin{array}[]{c}z_{t-1,1}\\ z_{t-1,2}\\ \epsilon_{t,1}\\ \epsilon_{t,2}\end{array}\end{bmatrix}\right).\end{gathered}$
@@
$\small p(z_{t-1,1},z_{t-1,2},z_{t,1},z_{t,2})=\frac{p(z_{t-1,1},z_{t-1,2},\epsilon_{t,1},\epsilon_{t,2})}{|\text{det }\mathbf{J}_{\mathbf{f}}|},$ (33)
@@
$\small\begin{gathered}\mathbf{J}_{\mathbf{f}}=\begin{bmatrix}\begin{array}[]{cccc}1&0&0&0\\ 0&1&0&0\\ \frac{\partial z_{t,1}}{\partial z_{t-1,1}}&\frac{\partial z_{t,1}}{\partial z_{t-1,2}}&\frac{\partial z_{t,1}}{\partial\epsilon_{t,1}}&0\\ \frac{\partial z_{t,2}}{\partial z_{t-1,1}}&\frac{\partial z_{t,2}}{\partial z_{t-1,2}}&\frac{\partial z_{t,2}}{\partial\epsilon_{t,1}}&\frac{\partial z_{t,2}}{\partial\epsilon_{t,2}}\end{array}\end{bmatrix}.\end{gathered}$
Given that this Jacobian is triangular, we can efficiently compute its determinant as $\prod_{i}\frac{\partial z_{t,i}}{\epsilon_{t,i}}$. Furthermore, because the noise terms are mutually independent, and hence $\epsilon_{t,i}\perp\epsilon_{t,j}$ for $j\neq i$ and $\epsilon_{t}\perp{\mathbf{z}}_{t-1}$, so we can with the RHS of Equation (33) as follows
$\small\begin{split}p(z_{t-1,1},z_{t-1,2},z_{t,1},z_{t,2})=p(z_{t-1,1},z_{t-1,2})\times\frac{p(\epsilon_{t,1},\epsilon_{t,2})}{|\mathbf{J}_{\mathbf{f}}|}=p(z_{t-1,1},z_{t-1,2})\times\frac{\prod_{i}p(\epsilon_{t,i})}{|\mathbf{J}_{\mathbf{f}}|}.\end{split}$ (34)
@@
$\small\begin{gathered}\underbrace{[\hat{{\mathbf{z}}}_{t-L},\cdots,{\hat{{\mathbf{z}}}}_{t-1},{\hat{{\mathbf{z}}}}_{t}]^{\top}}_{\mathbf{A}}\text{ mapped to }\underbrace{[{\hat{{\mathbf{z}}}}_{t-L},\cdots,{\hat{{\mathbf{z}}}}_{t-1},{\hat{\epsilon}}_{t,i}]^{\top}}_{\mathbf{B}},\text{ with }\mathbf{J}_{\mathbf{A}\rightarrow\mathbf{B}}=\begin{bmatrix}\begin{array}[]{cc}\mathbb{I}_{n_{s}\times L}&0\\ &\text{diag}\left(\frac{\partial r_{i,j}}{\partial{\hat{z}}_{t,j}}\right)\end{array}\end{bmatrix}.\end{gathered}$ (35)
Similar to Equation (34), we can obtain the joint distribution of the estimated dynamics subspace as:
$\log p(\mathbf{A})=\underbrace{\log p({\hat{{\mathbf{z}}}}_{t-L},\cdots,{\hat{{\mathbf{z}}}}_{t-1})+\sum^{n_{s}}_{i=1}\log p({\hat{\epsilon}}_{t,i})}_{\text{Because of mutually independent noise assumption}}+\log(|\text{det}(\mathbf{J}_{\mathbf{A}\rightarrow\mathbf{B}})|)$ (36)
@@
$\small\log p({\hat{{\mathbf{z}}}}_{t}|\{{\hat{{\mathbf{z}}}}_{t-\tau}\}_{\tau=1}^{L})=\sum_{i=1}^{n_{s}}p({\hat{\epsilon}_{t,i}})+\sum_{i=1}^{n_{s}}\log|\frac{\partial r_{i}}{\partial{\hat{z}}_{t,i}}|$ (37)
@@
A3.2 Evident Lower Bound
@@
$\begin{split}\ln p({\mathbf{x}}_{1:T})&=\ln\frac{p({\mathbf{x}}_{1:T},{\mathbf{z}}_{1:T})}{p({\mathbf{z}}_{1:T}|{\mathbf{x}}_{1:T})}=\mathbb{E}_{q({\mathbf{z}}_{1:T}|{\mathbf{x}}_{1:T})}\ln\frac{p({\mathbf{x}}_{1:T},{\mathbf{z}}_{1:T})q({\mathbf{z}}_{1:T}|{\mathbf{x}}_{1:T})}{p({\mathbf{z}}_{1:T}|{\mathbf{x}}_{1:T})q({\mathbf{z}}_{1:T}|{\mathbf{x}}_{1:T})}\\ &\geq\underbrace{\mathbb{E}_{q({\mathbf{z}}_{1:T}|{\mathbf{x}}_{1:T})}\ln p({\mathbf{x}}_{1:T|}|{\mathbf{z}}_{1:T})}_{L_{r}}-\underbrace{D_{KL}(q({\mathbf{z}}_{1:T}|{\mathbf{x}}_{1:T})||p({\mathbf{z}}_{1:T}))}_{L_{KLD}}=ELBO\end{split}$ (38)
A3.3 Model Details
We choose MICN [46] as the encoder backbone of our model on real-world datasets. Specifically, given the MICN extract the hidden feature, we apply a variational inference block and then a MLP-based decoder. Architecture details of the proposed method are shown in Table A1.
@@
Input:${\mathbf{x}}_{1:t}$ Observed time series Batch Size×$\times$×t×$\times$ ${\mathbf{x}}$ dimension
Dense |${\mathbf{x}}_{t}$| neurons Batch Size×$\times$×t×$\times$|${\mathbf{x}}_{t}$|
Concat zero concatenation Batch Size×$\times$×T×$\times$|${\mathbf{x}}_{t}$|
Dense n neurons Batch Size×$\times$×T×$\times$×n
@@
Input:${\mathbf{z}}_{1:T}$ Latent Variable Batch Size×$\times$×T×$\times$×n
Dense |${\mathbf{x}}_{t}$| neurons, Tanh Batch Size×$\times$×T×$\times$|${\mathbf{x}}_{t}$|
@@
Input: ${\mathbf{z}}_{1:T}$ Latent Variable Batch Size×$\times$(n+1)
Dense 128 neurons,LeakyReLU (n+1)$\times$128
Dense 128 neurons,LeakyReLU 128$\times$128
Dense 128 neurons,LeakyReLU 128$\times$128
Dense 1 neuron Batch Size×$\times$1
@@
Appendix A4 Experiment Details
A4.1 Simulation Experiment
A4.1.1 Data Generation Process
As for the temporally latent processes, we use MLPs with the activation function of LeakyReLU to model the sparse time-delayed and instantaneous relationships of temporally latent variables. Moreover, each independent noise $\epsilon_{t,i}$ is sampled from the distribution of normal distributions. We further let the data generation process from latent variables to observed variables be MLPs with the LeakyReLU units.
We provide 4 synthetic datasets. Dataset A has $5$ dimensions and a relatively dense latent causal process, which still satisfies the sparsity assumption. Dataset B has $3$ dimension, whose latent causal process is shown in Figure 4. Dataset C has $8$ dimensions, in which $z_{i,t}$ affects $z_{i,t+1}$ and $z_{i+1,t}$ (if applicable) only. Dataset Dense has $8$ dimensions as well, with a fully connected time-delayed transition and link instantaneous effects.
A4.1.2 Evaluation Metrics.
To evaluate the identifiability performance of our method under instantaneous dependencies, we employ the Mean Correlation Coefficient (MCC) between the ground-truth ${\mathbf{z}}_{t}$ and the estimated $\hat{{\mathbf{z}}}_{t}$. A higher MCC denotes a better identification performance the model can achieve. In addition, we also draw the estimated latent causal process to validate our method.
A4.1.3 More Simulation Experiment Results
We run each experiment 3 times, with seeds $769,770,771$. Variance results of the simulation datasets are shown in Table A2.
Table A2: Variance results on simulation data.
@@
Dense 0.037 0.005 0.001 0.021 0.015 0.002 0.009 0.002 0.039 0.016
A4.1.4 Implementation Details
For the implementation of baseline models, we utilized publicly released code for TDRL and iCRITIS. However, since the author did not release the code for G-CaRL, we implemented it ourselves based on the paper.
It is important to note that the original code of iCRITIS only accepts images as input. To adapt it to our needs, we replaced the encoder and decoder with a Variational Autoencoder, with the same hyperparameters used in IDOL.
A4.2 Experiment Details.
We use ADAM optimizer [102] in all experiments and report the mean squared error (MSE) and the mean absolute error (MAE) as evaluation metrics. All experiments are implemented by Pytorch on a single NVIDIA RTX A100 40GB GPU.
@@
A4.3 Real-world Dataset Description
@@
A4.4 More Experiments Results
@@
A4.5 Ablation Study
@@
To evaluate the effectiveness of individual loss terms, we also devise the two model variants. 1) IDOL-K: remove the KL divergence restriction of prior estiamtion. 2) IDOL-S: remove sparsity restriction of latent dynamics. Experiment results on the Jog dataset are shown in Figure A1. We can find that the sparsity restriction of latent dynamics plays an important role in the model performance, reflecting that the restriction of latent dynamics can benefit the identifiability of latent variables. We also discover that incorporating the KL divergence has a positive impact on the overall performance of the model, which shows the necessity of identifiability.
Appendix A5 Limitation
The proposed method can hardly be applied in the high-dimensional time series data. Therefore, how to relax this restriction and make it scalable to the high-dimensional real-world datasets would be a meaningful future direction.
Appendix A6 Broader Impacts
The proposed IDOL identifies the latent variables to create a transparent model with a certain degree of interpretability, thereby revealing the causal mechanism of general time series data.

### v3

@@ E.1 Prior Likelihood Derivation
[…]
@@
$p(z_{t-1,1},z_{t-1,2},z_{t,1},z_{t,2})=\frac{p(z_{t-1,1},z_{t-1,2},\epsilon_{t,1},\epsilon_{t,2})}{|\text{det }\mathbf{J}_{\mathbf{f}}|},$ (A30)
@@
[…]
Given that this Jacobian is triangular, we can efficiently compute its determinant as $\prod_{i}\frac{\partial z_{t,i}}{\epsilon_{t,i}}$. Furthermore, because the noise terms are mutually independent, and hence $\epsilon_{t,i}\perp\epsilon_{t,j}$ for $j\neq i$ and $\epsilon_{t}\perp{\mathbf{z}}_{t-1}$, so we can with the RHS of Equation (A30) as follows
$\begin{split}p(z_{t-1,1},z_{t-1,2},z_{t,1},z_{t,2})=p(z_{t-1,1},z_{t-1,2})\times\frac{p(\epsilon_{t,1},\epsilon_{t,2})}{|\mathbf{J}_{\mathbf{f}}|}=p(z_{t-1,1},z_{t-1,2})\times\frac{\prod_{i}p(\epsilon_{t,i})}{|\mathbf{J}_{\mathbf{f}}|}.\end{split}$ (A31)
@@
[…]
Similar to Equation (A31), we can obtain the joint distribution of the estimated dynamics subspace as:
$\log p(\mathbf{A})=\underbrace{\log p({\hat{{\mathbf{z}}}}_{t-L},\cdots,{\hat{{\mathbf{z}}}}_{t-1})+\sum^{n_{s}}_{i=1}\log p({\hat{\epsilon}}_{t,i})}_{\text{Because of mutually independent noise assumption}}+\log(|\text{det}(\mathbf{J}_{\mathbf{A}\rightarrow\mathbf{B}})|)$ (A33)
@@
$\log p({\hat{{\mathbf{z}}}}_{t}|\{{\hat{{\mathbf{z}}}}_{t-\tau}\}_{\tau=1}^{L})=\sum_{i=1}^{n_{s}}p({\hat{\epsilon}_{t,i}})+\sum_{i=1}^{n_{s}}\log|\frac{\partial r_{i}}{\partial{\hat{z}}_{t,i}}|$ (A34)
@@
E.2 Evident Lower Bound
@@
$\begin{split}\ln p({\mathbf{x}}_{1:T})&=\ln\frac{p({\mathbf{x}}_{1:T},{\mathbf{z}}_{1:T})}{p({\mathbf{z}}_{1:T}|{\mathbf{x}}_{1:T})}=\mathbb{E}_{q({\mathbf{z}}_{1:T}|{\mathbf{x}}_{1:T})}\ln\frac{p({\mathbf{x}}_{1:T},{\mathbf{z}}_{1:T})q({\mathbf{z}}_{1:T}|{\mathbf{x}}_{1:T})}{p({\mathbf{z}}_{1:T}|{\mathbf{x}}_{1:T})q({\mathbf{z}}_{1:T}|{\mathbf{x}}_{1:T})}\\ &\geq\underbrace{\mathbb{E}_{q({\mathbf{z}}_{1:T}|{\mathbf{x}}_{1:T})}\ln p({\mathbf{x}}_{1:T|}|{\mathbf{z}}_{1:T})}_{L_{r}}-\underbrace{D_{KL}(q({\mathbf{z}}_{1:T}|{\mathbf{x}}_{1:T})||p({\mathbf{z}}_{1:T}))}_{L_{KLD}}=ELBO\end{split}$ (A35)
As for the time-series forecasting task, we let ${\mathbf{x}}_{1:t}$ and ${\mathbf{x}}_{t+1:T}$ be the historical and future observed variables, then the ELBO can be further derived as follows:
$\begin{split}&\ln p({\mathbf{x}}_{t+1:T},{\mathbf{x}}_{1:t})=\ln\frac{p({\mathbf{x}}_{t+1:T},{\mathbf{z}}_{1:T},{\mathbf{x}}_{1:t})}{p({\mathbf{z}}_{1:T}|{\mathbf{x}}_{1:t},{\mathbf{x}}_{t+1:T})}=\ln\frac{p({\mathbf{x}}_{t+1:T},{\mathbf{z}}_{1:t},{\mathbf{z}}_{t+1:T},{\mathbf{x}}_{1:t})}{p({\mathbf{z}}_{1:T}|{\mathbf{x}}_{1:T})}\\ \\ \geq&\mathbb{E}_{q({\mathbf{z}}_{1:T}|{\mathbf{x}}_{1:t})}\ln\frac{p({\mathbf{x}}_{t+1:T}|{\mathbf{z}}_{t+1:T})p({\mathbf{x}}_{1:t}|{\mathbf{z}}_{1:t})p({\mathbf{z}}_{1:T})}{q({\mathbf{z}}_{1:T}|{\mathbf{x}}_{1:t})}\\ \\ =&\underbrace{\mathbb{E}_{q({\mathbf{z}}_{1:T}|{\mathbf{x}}_{1:t})}\ln p({\mathbf{x}}_{1:t}|{\mathbf{z}}_{1:T})}_{\mathcal{L}_{r}}+\underbrace{\mathbb{E}_{q({\mathbf{z}}_{1:T}|{\mathbf{x}}_{1:t})}\ln p({\mathbf{x}}_{t+1:T}|{\mathbf{z}}_{1:T})}_{\mathcal{L}_{pre}}-\underbrace{D_{KL}(q({\mathbf{z}}_{1:T}|{\mathbf{x}}_{1:t})||p({\mathbf{z}}_{1:T}))}_{\mathcal{L}_{KLD}},\end{split}$ (A36)
where $\mathcal{L}_{r}$ denotes the reconstruct loss of historical data, $\mathcal{L}_{pre}$ denotes the forecasting loss of future data.
E.3 Model Details
We choose MICN wang2022micn as the encoder backbone of our model on real-world datasets. Specifically, given the MICN extracts the hidden feature, we apply a variational inference block and then an MLP-based decoder. Architecture details of the proposed method are shown in Table A4.
E.3.1 Reproducibility of Simulation Experiments.
For the implementation of baseline models, we utilized publicly released code for TDRL and iCRITIS. However, since the author did not release the code for G-CaRL, we implemented it ourselves based on the paper. It is important to note that the original code of iCRITIS only accepts images as input. To adapt it to our needs, we replaced the encoder and decoder with a Variational Autoencoder, with the same hyperparameters used in IDOL. Our code is modified based on the code of TDRL, and shared hyperparameters remain the same.
E.3.2 Reproducibility of Real-world Experiments.
The model details of our method are shown in Table A4. For a fair comparison, we employ the official implementations and use the default hyperparameters. Since our model is modified from the official implementations of TDRL, both our model and the baselines share the same hyperparameters. To achieve the results from the well-fit baselines, we employed the default hyper-parameters and tried different values of learning rate for the best models. Please refer to Table A5 to A11 for the implementation details of the baselines.
@@ E.3 Model Details
[…]
@@
[…]
@@
[…]
Dense 128 neurons,LeakyReLU $\times128$
Dense 128 neurons,LeakyReLU $\times128$
[…]
@@
Table A5: TDRL architecture details. $T$: length of time series. $|x_{t}|$: input dimension. $n$: latent dimension. LeakyReLU: Leaky Rectified Linear Unit. Tanh: Hyperbolic tangent function.
Configuration Description Output
$\phi$ Latent Variable Encoder
[…]
$\psi$ Decoder
[…]
r Modular Prior Networks
[…]
Dense 128 neurons,LeakyReLU $\times128$
Dense 128 neurons,LeakyReLU $\times128$
[…]
JacobianCompute Compute log(det(J)) Batch Size
Table A6: CARD architecture details. $T$: length of time series. $|x_{t}|$: input dimension. $p$: patch number. $n$: patch length.
Configuration Description Output
[…]
Table A7: iTransformer architecture details. $T$: length of time series. $|x_{t}|$: input dimension.
Configuration Description Output
[…]
Table A8: Auoformer architecture details. $T$: length of time series. $|x_{t}|$: input dimension.
Configuration Description Output
[…]
Table A9: TimesNet architecture details. $T$: length of time series. $|x_{t}|$: input dimension.
Configuration Description Output
[…]
Table A10: MICN architecture details. $T$: length of time series. $|x_{t}|$: input dimension. LeakyReLU: Leaky Rectified Linear Unit. Tanh: Hyperbolic tangent function.
Configuration Description Output
[…]
Table A11: FITS architecture details. $T$: length of time series. $|x_{t}|$: input dimension. $p$: patch number. $n$: patch length.
Configuration Description Output
[…]
Appendix F Experiment Details
F.1 Simulation Experiment
F.1.1 Data Generation Process
As for the temporally latent processes, we use MLPs with the activation function of LeakyReLU to model the sparse time-delayed and instantaneous relationships of temporally latent variables. For all datasets, we set sequence length as $5$ and transition lag as $1$. That is:
$z_{t,i}=\left(LeakyReLU(W_{i,:}\cdot\mathbf{z}_{t-1},0.2)+V_{<i,i}\cdot\mathbf{z}_{t,<i}\right)\cdot\epsilon_{t,i}+\epsilon_{t,i},$
where $W_{i,:}$ is the $i$-th row of $W$ and $V_{<i,i}$ is the first i1i-1 columns in the $i$-th row of $V$. Moreover, each independent noise $\epsilon_{t,i}$ is sampled from the distribution of normal distribution. We further let the data generation process from latent variables to observed variables be MLPs with the LeakyReLU units.
We provide 6 synthetic datasets, with 3,5,8,8,8,16 latent variables from A to F. For dataset A, we have $W_{A}=[[1,1,0],[0,1,0],[0,0,1]]$. For dataset $B$, we have $W_{B}$ as an eye matrix with 2 extra nonzero entries. For dataset $C,F$, we have $W_{C},W_{F}$ as eye matrices. For datasets $D$ and $E$, $W_{D}$ and $W_{E}$ are dense matrices with all nonzero entries, which are generated by the data generator of TDRL(yao2022temporally). When it comes to $V$, we have and only have $V_{i-1,i}=1\ \ \forall i>0$ for dataset $A,B,C,E,F$. We also set $V_{D}=\mathbf{0}$, which means that there are no instantaneous effects.
The total size of the dataset is 100,000, with 1,024 samples designated as the validation set. The remaining samples are the training set.
F.1.2 Evaluation Metrics.
To evaluate the identifiability performance of our method under instantaneous dependencies, we employ the Mean Correlation Coefficient (MCC) between the ground-truth ${\mathbf{z}}_{t}$ and the estimated $\hat{{\mathbf{z}}}_{t}$. A higher MCC denotes a better identification performance the model can achieve. In addition, we also draw the estimated latent causal process to validate our method. Since the estimated transition function will be a transformation of the ground truth, we do not compare their exact values, but only the activated entries.
F.1.3 More Simulation Experiment Results
We run each experiment 3 times, with seeds $769,770,771$. Standard deviation results of the simulation datasets are shown in Table A12.
Table A12: Standard Deviation results on simulation data.
@@ F.1.3 More Simulation Experiment Results
D 0.006 0.003 0.051 0.010 0.028 0.074 0.034 0.027 0.033 0.007
E 0.037 0.005 0.001 0.021 0.015 0.002 0.009 0.002 0.039 0.016
F 0.029 0.018 0.004 - 0.165 0.134 0.014 0.031 0.018 0.007
F.1.4 Ablation Study
To further illustrate the significance of sparsity, we compare our IDOL model with its variant that lacks a sparsity constraint (IDOL-S). The experimental results on the synthetic dataset are presented in Table A13. The result shows that sparsity constraint is crucial to the model.
Table A13: MCC results of IDOL and IDOL-S with mean and standard deviation.
IDOL IDOL-S
MCC 0.8595 (0.0599) 0.8498 (0.0481)
@@ F.1.4 Ablation Study
F.2 Real-world Experiments
F.2.1 Real-world Dataset Description
@@
The sample number of different datasets are shown in Table A16.
Table A16: Sample number of different datasets.
Dataset Size
Humaneva Gestures 1686
Jog 2050
ThrowCatch 1021
Walking 2454
Human Discussion 13759
Greeting 8416
Purchases 7480
SittingDown 16438
Walking 16257
WalkTogether 10658
F.2.2 More Experiments Results
@@
F.2.3 Qualitative Results
To further show the effectiveness of the proposed method, we also randomly select a batch of test samples and visualize the forecasting results of different baselines. Visualization results in Figure A9 show the predicted results of “Discussion”, “Greeting”, and “Purchases”, respectively, which illustrate how the forecasting results align the ground-truth motions, where the red lines denote the ground-truth motions and the lines in other colors denote the forecasting results of different methods. According to the experiment results, we can find that the forecasting results of our method achieve the best alignment of the ground truth. In the meanwhile, the other methods that do not consider the instantaneous dependencies of latent variables may generate some exaggerated body movements, for example, the hyperflexed knee joint and the actions that violate the physics rules, which reflects the importance of considering the instantaneous dependencies of latent causal process of time series data.
Figure A9: The illustration of visualization of different methods of Walking. The red lines denote the ground-truth motions, those lines in other colors denote the prediction motions of different methods.
F.2.4 Ablation Study
@@
To evaluate the effectiveness of individual loss terms, we also devise the two model variants. 1) IDOL-K: remove the KL divergence restriction of prior estimation. 2) IDOL-S: remove sparsity restriction of latent dynamics. Experiment results on the Jog dataset are shown in Figure A8. We can find that the sparsity restriction of latent dynamics plays an important role in the model performance, reflecting that the restriction of latent dynamics can benefit the identifiability of latent variables. We also discover that incorporating the KL divergence has a positive impact on the overall performance of the model, which shows the necessity of identifiability.
Table A17: Mean and standard deviation of wall-clock training times for one epoch.
Methods IDOL TDRL CARD FITS MICN iTransformer TimesNet Autoformer
Second 65.960(7.556) 33.902(5.781) 62.994(3.035) 92.941(0.715) 45.547(13.034) 38.155(6.441) 324.941(25.286) 51.648(15.673)
Appendix G Complexity Analysis
G.1 Wall-clock Training Times
To evaluate the computational complexity of our method, we provide the wall-clock training times of different methods. Specifically, we use a consistent hardware setup including the same GPU, CPU, and memory configurations for each model to ensure comparability. Sequentially, in our codes, we wrap the training process in a timer to measure the actual elapsed time from start to finish of one epoch. For each method, we repeat three different times and further report the mean and standard deviation. The wall-clock training times are shown in Table A17.
Based on the experimental results, we draw the following conclusions:
  • Compared to our baseline model TDRL, the wall-clock training time of the proposed IDOL is nearly twice as slow. Theoretically speaking, the primary computational difference is that IDOL calculates the Jacobian matrix for both time-delayed and instantaneous relationships, while TDRL only calculates the time-delayed component.
  • Compared to other mainstream baselines, our IDOL method is slower than some models like MICN and iTransformer. However, our method is still faster than models like FITS.
G.2 Model Efficiency Analysis
To evaluate the model efficiency, we provide model efficiency comparison on the low-dimension dataset (e.g., the Human dataset) and the high-dimension dataset (the MoCap dataset) by evaluating the model efficiency from three aspects: forecasting performance, training speed, and memory footprint. As shown in Figure A10, in low-dimensional datasets, IDOL performs nearly as efficiently as top methods like Autoformer and MICN and outperforms others like CARD. However, in high-dimensional datasets (117 observations), IDOL requires more training time due to the added complexity of Jacobian calculations for instantaneous effects.
Figure A10: Model efficiency comparison for one training step on different datasets.
Appendix H More Discussions
H.1 Analysis of High-Dimensional Data
Identifying causal relationships in high-dimensional latent variable settings is a well-recognized challenge in the causal inference community (lopez2022large; cheng2024cuts+). As the dimensionality increases, the complexity of identifying causal structures grows due to the expanding search space. We conduct experiments on both synthetic and real-world datasets to verify it and propose a potential solution.
H.1.1 Simulation Experiments
As for the experiment on the synthetic datasets, we follow the same data generation process to generate simulation datasets with latent variable dimensions of 8, 16, 24, and 32. All these datasets share a similar latent causal process: chain-like instantaneous effects and ono-to-one temporal effects. We then measured the Mean Correlation Coefficient (MCC) between the ground truth ${\mathbf{z}}_{t}$ and the estimated $\hat{{\mathbf{z}}}_{t}$. The experimental results are presented in Table A18. According to the experiment result, as the dimension of latent variables increases, the value of MCC is still acceptable, despite possible performance loss in high-dimensional problems.
Table A18: Experiments results of IDOL on simulation data on different dimensions of latent variables.
Dimension 8 16 24 32
MCC 0.9801(0.002) 0.9747 (0.0029) 0.9243 (0.0173) 0.8640 (0.0071)
H.1.2 Real-world Experiments
For the real-world datasets, we consider the CMU-MoCap dataset [^3] http://mocap.cs.cmu.edu/. It contains various motion capture recordings and 117 skeleton-based measurements, which is a higher dimensionality compared to the Human dataset. We choose the Running and Soccer motions and take the input-prediction length as 50-25 and 50-50, respectively. As shown in Table A19, the IDOL model achieved a comparable forecasting performance in the high-dimensional dataset Running.
H.1.3 Potential Solution for High-Dimension Data
To better address this challenge, here we propose several potential solutions to more effectively address the challenges of high-dimensional time-series data. One idea is to make use of the divide-and-conquer strategy. One possible way is to leverage independent relations in the measure of time series data, if any. For instance, if processes $X_{1}:=\{x_{t,1}|t\in T\}$ and $X_{2}:=\{x_{t,2}|t\in T\}$ happen to be independent of processes $X_{3}$ and $X_{4}$, then we can just learn the underlying processes for $(X_{1},X_{2})$ and $(X_{3},X_{4})$ separately. Another potential way is to use the conditional independent relations in the measured time series data. For example, if processes $X_{1}$ and $X_{2}$ are independent from $X_{3}$ and $X_{4}$ given $X_{5}$ and $X_{6}$, then we can just learn the underlying processes for $(X_{1},X_{2},X_{5},X_{6})$ and $(X_{3},X_{4},X_{5},X_{6})$ separately. In this way, we can reduce the search space and further reduce the complexity even in high-dimensional time series data. We hope that some other developments in reducing the computational load in deep learning can also be helpful.
Table A19: Experiment results of the CMU MoCap datasets.
Motion
Predicted
Length
IDOL TDRL CARD FITS MICN iTransformer TimesNet Autoformer
MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
Running 50-25 0.110 0.082 0.448 0.108 0.998 0.135 0.076 0.064 0.658 0.135 0.458 0.106 2.616 0.202 1.033 0.179
50-50 0.286 0.109 1.779 0.167 2.217 0.153 0.266 0.103 0.978 0.161 1.831 0.170 4.720 0.252 3.944 0.283
Soccer 50-25 0.022 0.043 0.026 0.047 0.206 0.082 0.076 0.065 0.057 0.066 0.032 0.051 0.063 0.068 0.211 0.105
50-50 0.079 0.071 0.084 0.073 0.397 0.108 0.265 0.103 0.284 0.120 0.133 0.082 0.392 0.107 0.452 0.143
H.2 Noisy Environments
In some real-world scenarios, it is possible for the mixing process not to be invertible, for example, if the mixing process is highly noisy in each observed process, which might be the case in financial data (hu2008instrumental). There are some developments relying on the additive noise assumptions. If one makes strong assumptions on the noise, such as assuming an additive noise model, it is possible to develop a certain type of identifiability just like the extension of nonlinear ICA to the additive noise model case in (khemakhem2020variational). However, general approaches to deal with non-parametric noise terms to developed in this field in the future. To further evaluate the insight of this potential solution, we have conducted experiments on synthetic data. Specifically, we first follow Equation (2) to generate latent variables with temporal and instantaneous dependencies and generate observed variables by ${\mathbf{x}}_{t}={\mathbf{g}}({\mathbf{z}}_{t},\varepsilon)$, where $\varepsilon$ is the noise introduced to the mixing process. Experiment results in Table A20 show that our method can still achieve relatively good identifiability results even in a noisy environment, proving the potential of our insight.
Table A20: The mean and standard deviation of MCC of the IDOL model under noise environment with different noise scales. The noise scale is defined as the ratio of the noise variance to the observation variance, expressed as a percentage.
Noise Scale 0.1 0.3 0.5 0.7
MCC 0.9257 0.8362 0.7381 0.6095
H.3 Analysis of Sample Size
To assess the impact of sample size on identification performance, we conduct an ablation study focusing on sample size. Specifically, we create subsets of Dataset A containing 1,000, 10,000, 25,000, and 50,000 samples. For each dataset, we used the same hyperparameters, such as learning rate, random seed, and batch size. We repeated the experiment three times with different random seeds and reported the values of mean and standard deviation. The results in Table A21 show that as the sample size decreases, the performance of the model gradually declines. However, even with just 1k samples, our method achieves relatively good performance, demonstrating its robustness in small sample scenarios.
Table A21: The mean and standard deviation of MCC on subsets of dataset A with different sizes.
Sample Size 1,000 10,000 25000 50,000 100,000
MCC 0.853(0.102) 0.884(0.011) 0.912(0.025) 0.945(0.023) 0.965(0.029)
H.4 Subsampled Time Series
Sequentially, we consider another case where the time series data are sampled with low resolutions. When the time series data is sampled with low frequency, additional edges are introduced into the Markov network, making it denser. We can further assume a sparse mixing process zheng2022identifiability to achieve identifiability. Specifically, When conditioned on historical information that provides sufficient changes, the sparse mixing procedure assumption imposes structural constraints on the mapping from estimated to true latent variables. This compensates for potentially insufficient distribution changes, enabling identifiability even when the time series data with low-sample regimes. To further evaluate this insight, we also conduct experiments on synthetic datasets with low-sample regimes. Specifically, we first generate time series data with temporal and instantaneous dependencies. Then we randomly subsample the synthetic time series data. To enforce the sparsity of the estimated mixing process, we add an extra L1 constraint regarding the partial differential between the estimated observed and latent variables, i.e. $|\frac{\partial\hat{{\mathbf{x}}}_{t}}{\partial\hat{{\mathbf{z}}}_{t}}|_{1}$. Experiment results are shown in Table A22.
Table A22: The mean and standard deviation of MCC between the standard IDOL and the IDOL with sparse mixing constraint.
Model IDOL+Sparse Mixing Constraint IDOL
MCC 0.837(0.078) 0.786(0.085)
From the experimental results, we can draw the following conclusions: 1) When time series data is sampled at low frequency, the identifiable performance of the IDOL model decreases. This is because the causal process between latent variables becomes dense, which is consistent with our theoretical results. 2) by employing a sparse mixing constraint, we found that the identifiability performance of the model has been significantly improved, which proves that the above insight is effective.